OpenStudy (anonymous):

For which equation does the point (10,74) lie on the graph?

Nnesha (nnesha):

\(\huge\color{Green}{{\rm welcome}\rm~to~open~study!!!!}\) \(\Huge \color{blue}{\star^{ \star^{\star:)}}}\Huge \color{red}{\star^{ \star^{\star:)}}}\) \(\Huge \color{orange}{\star^{ \star^{\star:)}}}\Huge \color{purple}{\star^{ \star^{\star:)}}}\)

Nnesha (nnesha):

equations ???

OpenStudy (anonymous):

yea equations

Nnesha (nnesha):

to check it you have to replace x and y by (10,74) <--this order pair where first number is x and 2nd one is y and if you get both side equal then that is your answer

Nnesha (nnesha):

which equations ?? do u have answer choices or any equation

OpenStudy (anonymous):

yes

Nnesha (nnesha):

so replace x and y into that equations

OpenStudy (alexandervonhumboldt2):

so to do it substitute point corrdinates (x, y) into equation and check if it will be true

OpenStudy (alexandervonhumboldt2):

like examle equation y=3x-9 substitue 74=3*10-9 and check if it is coorrect

Nnesha (nnesha):

@Daina900 did you understand ?? :)

OpenStudy (anonymous):

the equations are D=2.4t+32 ,D=4.2t+32 ,D=6t+32

Nnesha (nnesha):

oka lets do first one together lets say D= y T= x

Nnesha (nnesha):

 replace variables by their values substitute 74 into D and substitute 10 into x